The perimeter of a rectangular garden is 96m. The length of the garden is 8 more than the width. Find the length and width of the garden.
Answer by BGutridge(15)

Perimeter = 96
let x represent the width. Then the length is represented by x+8
The equation to solve then would be 2(x+x+8)=96 which simplifies to 4x+16=96
Isolating the variable leaves 4x=80, or x=20
The length is 20m and the width is 28m. Checking...
2x20=40 + 2x28=56. 40+56=96